A ferry leaves a harbor and travels to a resort island at an average speed of 22 mph. On the return trip, the ferry travels at an average speed of 18 mph because of fog. The total time for the trip is 5 h. How far is the island from the harbor
Found 2 solutions by mananth, amoresroy:
Answer by mananth(16949)

A ferry leaves a harbor and travels to a resort island at an average speed of 22 mph. On the return trip, the ferry travels at an average speed of 18 mph because of fog. The total time for the trip is 5 h. How far is the island from the harbor:
..
let distance be x
forward speed = 22 mph
time = x/22
...
return speed = 18 mph
return time = x/18
..
x/18+x/22=5
(22x+18x) / 18*22 =5
40x=5*18*22
x=5*18*22 /40
distance = 49.5 miles
...
m.ananth@hotmail.ca

Answer by amoresroy(361)

Let T = travel time to island in hours
5-T = travel time back to harbor in hours
D = distance of island from harbor
= 22 T
22T = 18(5-T)
22T = 90 - 18T
40T = 90
T = 2.25
D = 2.25 (22)
D = 49.5 miles